36-726 Statistical Practice Homework 02

This week you will be working on ethics while I am working on your papers.

To-do list for this week:

For Today:

• Turn in your 3-4 page preliminary reports.

• Short class. Briefly discuss this assignment.

For Wednesday:

• We will read and discuss the following paper, on JSTOR.

Finney, David J. (1991) Ethical aspects of statistical practice. Biometrika, 47, 331–339.

It will be good to read this in conjunction with the 1989 ASA statistical guidelines at

http://www.tcnj.edu/~asaethic/asagui.html or the longer 1999 guidelines at
http://www.amstat.org/profession/index.cfm?fuseaction=ethicalstatistics.

• For discussion Wednesday, please the following people lead discussion^{1} of the following sections:

– Finney, sections 1, 2, 3, & the ASA guidelines listed above: BJ & Everyone.

– Finney, section 4: Drew, April, Zhanwu

– Finney, section 5: Joey, Gaia, Gary

– Finney, section 6: Abby, Nathaniel, Shuhei

– Finney, section 7: Steve, Yi, Emily

For Friday: No class. But please go to the following web site:

https://cme.hs.pitt.edu/servlet/IteachControllerServlet?actiontotake=displaymainpage

and select the link called Responsible Conduct of Research. Next select the link called Research Integrity. Create an account for yourself and start the module on research integrity.

Due Monday Feb 5: After successfully completing the module on research integrity, please save an electronic copy of your certificate (or print it out as a last resort) and email it to me. You should retain a copy for yourself for future reference.

1To lead the discussion, write down 2-3 major points in the section, and also a couple questions that you’d like to answer or have answered in class.